# CARE: a Benchmark Suite for the Classification and Retrieval of Enzymes

### v2

@@ 1 Introduction
Identifying the specific chemical reactions that an enzyme is capable of performing (i.e. the enzyme’s function) is a key first step for many applications [5]. While hundreds of millions of proteins have been sequenced, less than 1% are annotated with function [6]. For instance, during standard metagenomic analyses, enzyme genes are annotated for their functions, which enables determination of an organism’s critical metabolic pathways and specialization [7, 8, 9]. When applied to chemical synthesis, enzyme annotations are needed to identify catalysts that can replace existing step(s) in drug synthesis procedures–known as retrobiosynthesis [10]. Likewise, enzyme engineering for new-to-nature function [11] involves discovering an enzyme starting point with a desired function before improving activity for that function using protein engineering techniques [12, 13]. Historically, similarity search algorithms, most notably BLAST, have been the most common methods used to assign function to protein sequences [14, 15]. These methods work by finding similar sequences in annotated reference databases, as similar sequences are likely to share function. However, up to one third of genes from bacterial genomes cannot be assigned function using existing methods [7]. As such, there is a need for more abundant, high-quality annotations of enzymes and automated workflows to identify enzymes with desired functions.
@@ 4 Task 1: Enzyme Classification
Task 1 benchmarking results. Benchmarking results for Task 1 are summarized in Table 2. We start with two methods as baselines, classification using a random order of EC numbers (Random), and Diamond BLAST at the protein sequence level, herein referred to as BLASTp, which is a workhorse bioinformatics tool that performs local-alignment to determine the most similar sequence(s) given a target query and a database [14].
Table 2: Performance of various methods on Task 1. Performance is measured as k=1 classification accuracy (%). However, for the promiscuous split, we use k=(number of true ECs) for Random and CLEAN and report the average accuracy across all true EC numbers. For this split, we did not evaluate ChatGPT and Pika, as we prompted for only a single EC in the response. For BLAST we use all EC numbers associated to the top single sequence hit. More details are provided in Section A.4. CLEAN is a state-of-the-art method at the time of publication. Other methods such as ProteInfer, HiFi-NN, Enzhier, ProtEx, PhiGnet, etc. could also be benchmarked here.
@@
While there exist many ML tools to directly perform EC number classification, CLEAN [20] and several others [51, 9, 24, 52, 55] are a few that seem to report the current state-of-the-art with comparable performance. Some of the latter methods were not yet publicly available at the time of this study, and other recent retrieval methods would also be promising approaches to test, but we opted to focus on benchmarking on CLEAN here. CLEAN generally performs better than BLASTp across the board, but there are certain cases where the performance is similar. ChatGPT was also tested, but it appeared to often hallucinate as it was forced to provide an answer, with results similar to the random baseline. Pika seems to bridge the gap between standard LLMs and enzyme classification, but its performance is not as good as standard tools like BLASTp [66].
Overall, these results suggest that there is still room to improve classification of protein sequences with low sequence identity (<30%) and that lie near multiple activity peaks (Price). The promiscuous split is not easy for all methods, even though the test sequences have high sequence identity to the train set – CLEAN sometimes misses EC numbers. Interestingly, BLAST performs close to the state-of-the-art, even when compared to more complex ML models. Future models could take advantage of this finding to augment training, as ProtEx does [24]. Additional details on the implementation of each method can be found in Appendix A.4.
@@ 6 Discussion
We made an important design choice in the CARE benchmarks: to perform classification and retrieval at the coarse-grained level of EC numbers, which could be a limitation. Enzymes can often perform many reactions, meaning it is an acceptable assumption that enzymes belonging to the same EC number will share the capacity to perform similar reactions, even if they are not directly annotated for all reactions. The ultimate task in this domain is to perform direct reaction to protein sequence retrieval and vice-versa (as in CLIPZyme [78] and Reactzyme [96]). However, currently the data for validation is limited. As experimentalists obtain higher resolution annotations of proteins and their associated reactions, this will become more realistic. Some EC numbers are also incorrectly annotated, which are discussed in more detail here [61].
@@
There is also significant opportunity to use other modalities such as textual description and protein structure for more effective representations of enzymes and reactions, or to improve the text-based annotations. The textual annotations used in this work are direct textual references to the EC class, meaning that the models in task 2 that incorporate text (e.g. CREEP), are effectively learning the relationship between the EC as a word and the EC as a number, in addition to the similarity of the reaction. Other tools such as Pika which use multiple types of textual description extend this to learn relations about the enzyme, and could be used in future iterations of CREEP. Future work will involve incorporating structure and graph based representations [98] into CREEP, similar to those used in CLIPZyme [78]. We also plan to do a more detailed analysis of the representations learned by CREEP. The addition of textual description in CREEP potentially introduces indirect data leakage between the train and test sets, so future iterations of CARE will need to consider this. Future work should also consider how to include protein function prediction models that go beyond enzymes–or models that would like to use additional data/modalities [57, 99]–into the CARE evaluation framework. LAB-Bench provides a framework for evaluating scientific reasoning using language, which will increasingly intersect with CARE [72]. A major bottleneck we encountered was that many models, including language models, were not available or difficult to use, limiting our ability to include them as benchmarks. There is also room to improve the prompt engineering of language models which could be further explored to enhance the performance of these models.

### v3

@@ 1 Introduction
Identifying the specific chemical reactions that an enzyme is capable of performing (i.e. the enzyme’s function) is a key first step for many applications [5]. While hundreds of millions of proteins have been sequenced, less than 1% are annotated with function [6]. For instance, during standard metagenomic analyses, enzyme genes are annotated for their functions, which enables determination of an organism’s critical metabolic pathways and specialization [7, 8, 9]. When applied to chemical synthesis, enzyme annotations are needed to identify catalysts that can replace existing step(s) in drug synthesis procedures–known as retrobiosynthesis [10]. Likewise, enzyme engineering for new-to-nature function [11] involves discovering an enzyme starting point with a desired function before improving activity for that function using protein engineering techniques [12, 13]. Historically, similarity search algorithms, most notably BLAST, have been the most common methods used to assign function to protein sequences [14, 15]. These methods work by finding similar sequences in annotated reference databases, as similar sequences are likely to share function. More recently, searching based on structural similarity (Foldseek) has also demonstrated utility for assigning function [16, 17]. However, up to one third of genes from bacterial genomes cannot be assigned function using existing methods [7]. As such, there is a need for more abundant, high-quality annotations of enzymes and automated workflows to identify enzymes with desired functions.
@@ 4 Task 1: Enzyme Classification
Task 1 benchmarking results. Benchmarking results for Task 1 are summarized in Table 2. We start with two methods as baselines, classification using a random order of EC numbers (Random), and Diamond BLAST at the protein sequence level, herein referred to as BLASTp, which is a workhorse bioinformatics tool that performs local-alignment to determine the most similar sequence(s) given a target query and a database [14]. Additionally, we include search by structural similarity using Foldseek [16].
Table 2: Performance of various methods on Task 1. Performance is measured as k=1 classification accuracy (%). However, for the promiscuous split, we use k=(number of true ECs) for Random and CLEAN and report the average accuracy across all true EC numbers. For this split, we did not evaluate ChatGPT and Pika, as we prompted for only a single EC in the response. For BLAST and Foldseek, we use all EC numbers associated to the top single hit. More details are provided in Section A.4. CLEAN is a state-of-the-art method at the time of publication. Other methods such as ProteInfer, HiFi-NN, Enzhier, ProtEx, PhiGnet, etc. could also be benchmarked here.
@@
While there exist many ML tools to directly perform EC number classification, CLEAN [22] and several others [53, 9, 26, 54, 57] are a few that seem to report the current state-of-the-art with comparable performance. Some of the latter methods were not yet publicly available at the time of this study, and other recent retrieval methods would also be promising approaches to test, but we opted to focus on benchmarking on CLEAN here. CLEAN generally performs the best for enzymes with low sequence identity to known sequences, but the BLASTp and Foldseek baselines perform similarly. ChatGPT was also tested, but it appeared to often hallucinate as it was forced to provide an answer, with results similar to the random baseline. Pika seems to bridge the gap between standard LLMs and enzyme classification, but its performance is not as good as standard tools like BLASTp [67].
Overall, these results suggest that there is still room to improve classification of protein sequences with low sequence identity (<30%) and that lie near multiple activity peaks (Price). In fact, Foldseek has the best performance on the Price split, suggesting that structure can be conserved, even in scenarios where sequence is misleading. The promiscuous split is not easy for all methods; even though the test sequences have high sequence identity to the train set, CLEAN sometimes misses EC numbers. Interestingly, BLASTp and Foldseek perform close to the state-of-the-art, even when compared to more complex ML models. Future models could take advantage of this finding to augment training, as ProtEx does [26]. Additional details on the implementation of each method can be found in Appendix A.4.
@@ 6 Discussion
We made an important design choice in the CARE benchmarks: to perform classification and retrieval at the coarse-grained level of EC numbers, which could be a limitation. Enzymes can often perform many reactions, meaning it is an acceptable assumption that enzymes belonging to the same EC number will share the capacity to perform similar reactions, even if they are not directly annotated for all reactions. The ultimate task in this domain is to perform direct reaction to protein sequence retrieval and vice-versa (as in CLIPZyme [79] and Reactzyme [97]). However, currently the data for validation is limited, and we believe that many negative examples in those datasets may actually be feasible protein-reaction pairs. As experimentalists obtain higher resolution annotations of proteins and their associated reactions, this will become more realistic. Furthermore, our goal was to provide benchmarks for a broad range of functions, but model evaluation could also be performed on more specific classes of enzymes with direct protein-reaction pairs, as explored by EnzymeCAGE [98]. Other ways to hold out proteins with difficult-to-predict functions could also be explored. Some EC numbers are also incorrectly annotated, which are discussed in more detail here [62].
@@
There is also significant opportunity to use other modalities such as textual description and protein structure for more effective representations of enzymes and reactions, or to improve the text-based annotations. Gene context is also useful for annotating protein function and could be incorporated into future benchmarks [100]. The textual annotations used in this work are direct textual references to the EC class, meaning that the models in task 2 that incorporate text (e.g. CREEP), are effectively learning the relationship between the EC as a word and the EC as a number, in addition to the similarity of the reaction. Other tools such as Pika which use multiple types of textual description extend this to learn relations about the enzyme, and could be used in future iterations of CREEP. Future work will involve incorporating structure and graph based representations [101] into CREEP, similar to those used in CLIPZyme [79]. We also plan to do a more detailed analysis of the representations learned by CREEP. The addition of textual description in CREEP potentially introduces indirect data leakage between the train and test sets, so future iterations of CARE will need to consider this. Future work should also consider how to include protein function prediction models that go beyond enzymes–or models that would like to use additional data/modalities [59, 102]–into the CARE evaluation framework. LAB-Bench provides a framework for evaluating scientific reasoning using language, which will increasingly intersect with CARE [73]. A major bottleneck we encountered was that many models, including language models, were not available or difficult to use, limiting our ability to include them as benchmarks. There is also room to improve the prompt engineering of language models which could be further explored to enhance the performance of these models.
@@ A.4 Implementation Details for Benchmarking Methods
Foldseek. Structures for entries in the test sets were obtained from the AF2 database. Missing structures were folded with ESMFold, but a few rare cases ( 5) could not be successfully folded. Structures for the Price dataset were obtained with ESMFold. Structures for sequences in the training set were obtained from the AF2 database. While a small fraction of structures in the training set could not be retrieved from the AlphaFold database, we decided to proceed without these. Structure search was performed with Foldseek with default parameters, and only the top hit was used to assign function.